a community center sells 304 tickets for a b-ball game adult tickets are $4.00 and student tickets are $1.00. They collected $589.00 in sells. How many of eack tickets were sold?
Answer by Cromlix(4381)

Adult tickets = x
Student tickets = y
x + y = 304...1
4x + y = 589...2
Subtract (1) from (2)
3x = 285
x = 95
Substitute x = 95 in (1)
x + y = 304
95 + y = 304
y = 304 - 95
y = 209
Adult tickets sold = 95
Student tickets sold = 209.
Hope this helps.
:-)